OpenStudy (anonymous):

A culture started with 1000 bacteria. After 6 hours, it grew to be 1100 bacteria. Predict how many bacteria will be present after 15 hours. Round your answer to the nearest whole number.

OpenStudy (anonymous):

2500

OpenStudy (anonymous):

Wrong, 15/6=2.5 2.5*1100=2500 2500+1000=3500 3500

OpenStudy (anonymous):

you forgot to add the starting 1000